OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 36 inches. Find the width of the rectangle.

OpenStudy (anonymous):

perimeter is the distance around the rectangle, which would be l + w + l + w, or 2l + 2w. So now you have 2 equations: 1. l = w + 4 (length is 4 inches more than width) 2. 2l + 2w = 36 (perimeter is 36 inches) you can substitute "w + 4" for "l" in the second equation, because the first equation tells us they are equal: 2(w + 4) + 2w = 36 2w + 8 + 2w = 36 4w + 8 = 36 4w = 28 w = 7 Then substitute 7 for w in equation 1: l = 7 + 4 l = 11 Then check that we got the right answer. l = w + 4 11 = 7 + 4 11 = 11 (check!) 2l + 2w = 36 2(11) + 2(7) = 36 22 + 14 = 36 36 = 36 (check!)